# Self-similar decomposition of the hierarchical merger tree of dark matter halos

### v2

@@ 3.3 Dependence on Cosmology
These findings are consistent with expectations from the Extended Press-Schechter (EPS) formalism. To see this, we follow Dong et al. (2022) to calculate the specific merger rate using the EPS theory (their Equation (11)) and then convert the result to the PMF (their Equation (4)). We use the Eisenstein & Hu (1998) transfer function to calculate the linear power spectrum and adopt a top-hat filter in Fourier space to compute the variance of fluctuation on a given scale. The predictions for the PMFs under different cosmological parameters are presented in Figure 4.
In the left panel, the predicted PMFs are nearly universal for $\mu$ values greater than $10^{-3}$, regardless of variations in $\Omega_{M_{0}}$ and $\sigma_{8}$. The slopes of the predicted PMFs show slight differences below $10^{-3}$, where larger $\Omega_{M_{0}}$ values result in steeper slopes. This cosmology dependence also suggests a weak redshift dependence in a given cosmology, due to the evolution of the matter density and fluctuation amplitude with redshift. Verifying this residual dependence requires higher resolution simulations than those used in the current work.
Despite the detailed dependencies, we emphasize that significant variations in the PMFs can only be observed with very large variations in cosmological parameters (e.g., by a factor of two). For small (e.g., $\sim 10\%$) variations in the cosmological parameters, the variations in the PMFs will be negligible, and it remains a very good approximation to describe the PMFs as universal.
@@ 4.1 Accretion Level Distribution
These results have important implications for studying the evolution of subhalos. In particular, low-mass subhalos are dominated by high-level ones which are subject to "pre-processing" within their previous hosts. These objects are also likely still orbiting inside their host subhalos rather than the central subhalo in the final host. To model their evolution, it is thus more important to study their orbits around, and the gravitational influence from their host subhalos. The evolution of high-level subhalos could also be more complex due to multibody interactions among the subhalo group in addition to their interactions with the host (sub)halo. For example, Ludlow et al. (2009) found that some subhalos can be ejected far outside the host halo due to three-body interactions.
@@ 4.2 Subhalo Accretion Rate at Each Level
$=-g^{\prime}(\mu_{\rm acc})\mu_{\rm acc}\delta\ln M_{h}\mathrm{d}\ln m$
$=-g^{\prime}(\mu_{\rm acc})\delta\ln M_{h}\mathrm{d}\mu_{\rm acc},$
@@
It describes the accretion rate of level $\ell$ subhalos, originating from level $\ell-1$ subhalos, onto the host at a given redshift. Note $\mu_{\rm acc}$ is the mass ratio relative to the host mass at the redshift of the accretion. When $\ell=1$, the accretion rate becomes equivalent to the merger rate of progenitor halos.
@@
The predicted specific accretion rate at each level and their contributions to the total merger rate are shown in the left and right panels of Figure 6, respectively. In the left panel, the level-1 accretion rate is also compared against that of Dong et al. (2022) for the halo merger rate. [^2] The merger ratio in Dong et al. (2022), $\mu_{\rm Dong}$, is defined as the ratio between the progenitor masses prior to their merger, while in our case the ratio $\mu_{\rm acc}$ is between the subhalo peak mass and the host halo mass after the merger. We correct for the difference through $\mu_{\rm acc}=\frac{\mu_{\rm Dong}}{1+\mu_{\rm Dong}}$. Note both works are based on simulations that extend to $\mu_{\rm acc}\sim 5\times 10^{-4}$. Above this mass ratio limit, the two models are largely consistent with each other. The slopes of the two models differ more significantly at the low mass end. This can be attributed to the different approaches in deriving the models. In Dong et al. (2022), the halo merger rate is measured directly from the halo merger tree under certain definitions for halo merger (e.g., in time and distance threshold). Some post processing of the tree is also required to account for fly-by and splash-back events. In addition, the virial masses prior to the merger time are adopted to define the merger ratio, in contrast to the peak mass and host bound mass used in our work. These operations could lead to systematic differences in the merger rate from that obtained in this work.
In the right panel, the specific merger rate $f_{\ell}$ at each level has been normalized by the total rate, $f_{t}$, to show their relative contributions. The partitioning is similar to that shown in Figure 5. The level-1 mergers still dominate the high $\mu_{\rm acc}$ region. With a decrease in $\mu_{\rm acc}$, contributions of higher-level mergers increase. At $\mu_{\rm acc}=0.017$, the specific merger rates for the first two levels, $f_{1}$ and $f_{2}$, equalize. It occurs at a higher ratio compared with the transition point between $g_{1}$ and $g_{2}$.
@@ 4.3 Initial Merger Ratio Distribution
We note that some subhalos are observed with an excessively high ratio of $\xi>1$ in Figure 7, even though they constitute only a very small fraction (1%–4%) of the investigated population. The existence of these objects can be understood as due to systematics in building the merger tree. hbt+does not simply choose the most massive progenitor as the central subhalo during the merger, but relies on a kinematic distance to the host center to make the decision. Thus, a satellite may have a peak mass larger than that of the central. Furthermore, approximately 50% of these high-merger-ratio subhalos have experienced central-satellite switching during their evolutionary history. During this switching process, they may acquire peak masses comparable to or even exceeding those of central subhalos. After all, the two progenitors are close in mass in a $\xi\sim 1$ merger, so it is always somewhat arbitrary to select one as central. We thus refrain from further discussing and modeling the distribution in this part.
Figure 7: The probability distributions of the initial merger ratio, $\xi$, for subhalos of a given final mass ratio, $\mu$ (in different panels), and final subhalo level, $\ell$ (in different colors). The data points are measurements in the L600 simulation, while the solid lines show the model predictions. The vertical dashed curve highlights the $\mu$ value in each panel.
@@ 4.4 Accretion Redshift Distribution
With universal models available in the literature to predict the mass accretion history, $\tilde{M}(z)$, Equation (16) can be evaluated explicitly. For example, Zhao et al. (2009) developed a universal model for the median mass assembly history of dark matter halos with a given final mass at $z=z_{0}$ (see also van den Bosch, 2002; Wechsler et al., 2002; Li et al., 2007; Ludlow et al., 2014). With this specific model, we can construct the accretion history of high-level subhalos analytically.
@@
Figure 9: The accretion redshift distributions for subhalos of a given mass ratio, $\mu_{0}$, and host halo mass, $M_{0}$. The data points show the simulation measurements in L600, while the solid lines show the model predictions. Different colors show the results for different subhalo levels. The top and bottom panels show the results for $M_{0}=10^{13}\mathrm{M}_{\odot}\;h^{-1}$ and $M_{0}=10^{14}\mathrm{M}_{\odot}\;h^{-1}$, respectively, while the different columns show the results for different $\mu_{0}$ as labeled. The error bars show the Poisson noise.
@@ 5.1 Influences of Mass Definitions and Membership
We have tried various combinations of the different choices. Overall, the resulting mass functions can all be described by the double Schechter form. However, the level of universality varies depending on the mass and membership definitions. The optimal combination yielding the highest universality is the one adopted for our main results above, defining the peak mass as the maximum bound mass, the host halo mass as the sum of the subhalo bound masses, and the membership of subhalos according to the FoF catalog.
In Figure 10, we show examples of two other combinations. In the top panels, the host halo mass is defined as $M_{200c}$, and subhalos are counted only if they are within the corresponding virial radius, $R_{200c}$, of the host halo. Disrupted subhalos are also counted with their locations traced by their most bound particles. The peak mass is defined as the maximum bound mass as in the fiducial combination. In the bottom panels, an alternative virial definition, $M_{\rm vir}$, is adopted. In both cases, a larger mass and redshift dependence at the high $\mu$ end can be observed compared with our fiducial combination. At a fixed host mass bin, the PMF tends to decrease with the increase in redshift at the high $\mu$ end, while the low $\mu$ distribution remains universal. This is equivalently observed as a decrease in the shoulder scale of the exponential tail of the PMF towards high redshifts. The evolution is stronger in more massive hosts and weaker in less massive hosts, leading to larger discrepancies among the PMFs at high redshifts.
Figure 10: The PMFs obtained from L600 with host halo mass defined according to the virial definitions. Only subhalos (including disrupted ones) residing within the corresponding virial radius are counted. The top and bottom rows show results following the $M_{200c}$ and $M_{\rm vir}$ virial definitions, respectively. Dots of different colors show results in different halo mass bins. The black solid line in each panel is the model $g_{1}$ converted from the best fitting $g_{t}$ in Han et al. (2018) according to the recurrence relation. The gray-shaded region represents ±$\pm 10\%$ deviations from the model. The measurements from L600 have been corrected to the mass resolution of the Millennium-II simulation in order to compare against the Han et al. (2018) model (see section 5.2).
@@
These results are not difficult to understand. When two halos merge, the resulting FoF halo largely conserves the mass from both progenitors. Thus, when using the total bound mass of the FoF as a halo mass proxy, the increment in the halo mass can have a good scaling with the increment in progenitor mass. On the other hand, the spherical overdensity mass definitions do not respect the morphology and extent of the merging halos, so that the scaling between the progenitor mass and halo mass is more complex. This complexity can also be understood as an inconsistency between the boundary of the host halo and the orbital extent of its member particles and subhalos. Recent halo boundary definitions that incorporate the splashback or growth envelope of a halo, e.g., the depletion radius (Fong & Han, 2021; Gao et al., 2023), could potentially provide more self-consistent boundaries for the study of halo mergers and the resulting mass and subhalo distribution (Zhou & Han, 2023, 2025; Li & Han, 2021; Fong et al., 2022).
@@ 5.2 Correcting Excessive Mass from Unresolved Subhalos
Figure 12: The final subhalo mass functions in L600 at $z=0$ after being rescaled by $(M_{h}/10^{10}\mathrm{M}_{\odot}\;h^{-1})^{0.1}$. Dots of different colors show the results obatined from different halo mass bins. The gray solid line indicates the best fit (Equation (18)).
Figure 13: Left: Comparisons of $g_{1}$ obtained from the Kanli CDM simulation and L600 without peak mass correction. All of them have been normalized by the best-fitting formula in Equation (9). The solid and dashed lines represent results from L600 and Kanli CDM, respectively. In the plots, the PMFs are measured from two distinct host mass bins, $[10^{12},10^{13}]{\rm M_{\odot}}\,h^{-1}$ (blue points) and $[10^{13},10^{14}]{\rm M_{\odot}}\,h^{-1}$ (red points), in each simulation. Shaded regions indicate deviations of 5% from the fitting formula. Right: The comparisons of $g_{1}$ measured from two simulations after applying the mass correction to L600.
@@ 6 Conclusion
In this work, we have carried out a detailed analysis on the hierarchical origin of dark matter subhalos, using a set of N-body simulations. Applying the hbt+subhalo finder to the simulations, we classify subhalos into different levels according to their merger histories, and use the subhalo peak mass function (PMF) of each level to summarize their hierarchical origins, in which the peak mass of a subhalo is used as a proxy for the progenitor mass. Our main findings are summarized below.
  • The subhalo peak mass function (PMF) is very close to a universal function across halo mass and redshift. A single double-Schechter fit can describe the level-1 subhalo PMF to $10\%$ accuracy down to a peak-to-host mass ratio of $\mu=5\times 10^{-4}$, for the halo mass range of $10^{11}\mathrm{M}_{\odot}\;h^{-1}-10^{15}\mathrm{M}_{\odot}\;h^{-1}$ and redshift range of $z=0-3$ covered in our analysis.

### v3

@@ 3.3 Dependence on Cosmology
The results found here are consistent with analytical expectations in the EPS theory. In Appendix B, we evaluate the PMFs in EPS under varying cosmological parameters, and found that the PMFs are also nearly invariant under variations of $\Omega_{\rm M_{0}}$ and $\sigma_{8}$, but show a weak dependence on the $n_{s}$ parameter. Despite this, significant variations in the PMFs can only be observed with very large variations in the $n_{s}$ parameter (e.g., by a factor of two). For small variations (e.g., $\sim 10\%$) in the cosmology, it remains a very good approximation to describe the PMFs as universal.
@@ 4.1 Accretion Level Distribution
These results have important implications for studying the evolution of subhalos. In particular, low-mass subhalos are dominated by high-level ones which are subject to “pre-processing” within their previous hosts. These objects are also likely still orbiting inside their host subhalos rather than the central subhalo in the final host. To model their evolution, it is thus more important to study their orbits around, and the gravitational influence from their host subhalos. The evolution of high-level subhalos could also be more complex due to multibody interactions among the subhalo group in addition to their interactions with the host (sub)halo. For example, Ludlow et al. (2009) found that some subhalos can be ejected far outside the host halo due to three-body interactions.
@@ 4.2 Subhalo Accretion Rate at Each Level
$=-\mu_{\rm acc}g^{\prime}(\mu_{\rm acc})\delta\ln M_{h}\mathrm{d}\ln m$
$=-g^{\prime}(\mu_{\rm acc})\delta\ln M_{h}\mathrm{d}\mu,$
@@
It describes the accretion rate of level $\ell$ subhalos, originating from level $\ell-1$ subhalos, onto the host at a given redshift. When $\ell=1$, the accretion rate becomes equivalent to the merger rate of progenitor halos.
@@
The predicted specific accretion rate at each level and their contributions to the total merger rate are shown in the left and right panels of Figure 5, respectively. In the left panel, the level-1 accretion rate is also compared against that of Dong et al. (2022) for the halo merger rate. [^2] The merger ratio in Dong et al. (2022), $\mu_{\rm Dong}$, is defined as the ratio between the progenitor masses prior to their merger, while in our case the ratio $\mu_{\rm acc}$ is between the subhalo peak mass and the host halo mass after the merger. We correct for the difference through $\mu_{\rm acc}=\frac{\mu_{\rm Dong}}{1+\mu_{\rm Dong}}$. Note both works are based on simulations that extend to $\mu_{\rm acc}\sim 5\times 10^{-4}$. Above this mass ratio limit, the two models are largely consistent with each other. The slopes of the two models differ more significantly at the low mass end. This can be attributed to the different approaches in deriving the models. In Dong et al. (2022), the halo merger rate is measured directly from the halo merger tree under certain definitions for halo merger (e.g., in time and distance thresholds). Some post-processing of the tree is also required to account for fly-by and splash-back events. In our work, however, these events are automatically accounted for through the variation of the PMF at different time. In addition, the virial masses prior to the merger time are adopted to define the merger ratio, in contrast to the peak mass and host bound mass used in our work. These operations could lead to systematic differences in the merger rate from that obtained in this work.
In the right panel, the specific merger rate $f_{\ell}$ at each level has been normalized by the total rate, $f_{t}$, to show their relative contributions. The partitioning is similar to that shown in Figure 4. The level-1 mergers still dominate the high $\mu$ region. With a decrease in $\mu$, contributions of higher-level mergers increase. At $\mu_{\rm acc}=0.017$, the specific merger rates for the first two levels, $f_{1}$ and $f_{2}$, equalize. It occurs at a higher ratio compared with the transition point between $g_{1}$ and $g_{2}$.
@@ 4.3 Initial Merger Ratio Distribution
We note that some subhalos are observed with an excessively high ratio of $\xi>1$ in Figure 6, even though they constitute only a very small fraction (1%–4%) of the investigated population. The existence of these objects can be understood as due to systematics in building the merger tree. hbt+does not simply choose the most massive progenitor as the central subhalo during the merger, but relies on a kinematic distance to the center of mass of the host halo to make the decision. Thus, a satellite may have a peak mass larger than that of the central. Furthermore, approximately 50% of these high-merger-ratio subhalos have experienced central-satellite switching during their evolutionary history. During this switching process, they may acquire peak masses comparable to or even exceeding those of central subhalos. After all, the two progenitors are close in mass in a $\xi\sim 1$ merger, so it is always somewhat arbitrary to select one as central. We thus refrain from further discussing and modeling the distribution in this part.
Figure 6: The probability distributions of the initial merger ratio, $\xi$, for subhalos of a given final mass ratio, $\mu$ (in different panels), and final subhalo level, $\ell$ (in different colors). The data points are measurements in the L600 simulation, while the solid lines show the model predictions. The vertical dashed curve highlights the $\mu$ value in each panel. The errorbars show the Poisson noise estimated from the total subhalo count in each initial merger ratio bin.
@@ 4.4 Accretion Redshift Distribution
With universal models available in the literature to predict the mass accretion history, $\tilde{M}(z)$, Equation (16) can be evaluated explicitly. For example, Zhao et al. (2009) developed a universal model for the median mass assembly history of dark matter halos with a given final mass at $z=z_{0}$ (see also van den Bosch, 2002; Wechsler et al., 2002; Li et al., 2007; Ludlow et al., 2014; Liu et al., 2024). With this specific model, we can construct the accretion history of high-level subhalos analytically.
@@
Figure 8: The accretion redshift distributions for subhalos of a given mass ratio, $\mu_{0}$, and host halo mass, $M_{0}$. The data points show the simulation measurements in L600, while the solid lines show the model predictions. Different colors show the results for different subhalo levels. The top and bottom panels show the results for $M_{0}=10^{13}\mathrm{M}_{\odot}\;h^{-1}$ and $M_{0}=10^{14}\mathrm{M}_{\odot}\;h^{-1}$, respectively, while the different columns show the results for different $\mu_{0}$ as labeled. The errorbars show the Poisson noise estimated from the subhalo count in each redshift bin.
@@ 5.1 Influences of Mass Definitions and Membership
We have tried various combinations of the different choices. Overall, the resulting mass functions can all be described by the double Schechter form. However, the level of universality varies depending on the mass and membership definitions. The optimal combination yielding the highest universality is the one adopted for our main results above, defining the peak mass as the maximum bound mass, the host halo mass as the sum of the subhalo bound masses, and the membership of subhalos according to the FoF catalog. We have also tried using the FoF mass as the host halo mass, which performs better than the virial masses while the universality is slightly worse than that using our fiducial definition.
In Figure 9, we show examples of two other combinations. In the top panels, the host halo mass is defined as $M_{200c}$, and subhalos are counted only if they are within the corresponding virial radius, $R_{200c}$, of the host halo. Disrupted subhalos are also counted with their locations traced by their most bound particles. The peak mass is defined as the maximum bound mass as in the fiducial combination. In the bottom panels, an alternative virial definition, $M_{\rm vir}$, is adopted. In both cases, a stronger mass and redshift dependence at the high $\mu$ end can be observed compared with our fiducial combination. At a fixed host mass bin, the PMF tends to decrease with the increase in redshift at the high $\mu$ end, while the low $\mu$ distribution remains universal. This is equivalently observed as a decrease in the shoulder scale of the exponential tail of the PMF towards high redshifts. The evolution is stronger in more massive hosts and weaker in less massive hosts, leading to larger discrepancies among the PMFs at high redshifts.
Figure 9: The PMFs obtained from L600 with host halo mass defined according to the virial definitions. Only subhalos (including disrupted ones) residing within the corresponding virial radius are counted. The top and bottom rows show results following the $M_{200c}$ and $M_{\rm vir}$ virial definitions, respectively. The errorbars show the uncertainty on the average PMF estimated from the Poisson noise in the total subhalo count in each bin. Dots of different colors show results in different halo mass bins. The black solid line in each panel is the model $g_{1}$ converted from the best fitting $g_{t}$ in Han et al. (2018) according to the recurrence relation. The gray-shaded region represents ±$\pm 10\%$ deviations from the model. The measurements from L600 have been corrected to the mass resolution of the Millennium-II simulation in order to compare against the Han et al. (2018) model (see section 5.2).
@@
These results are not difficult to understand. When two halos merge, the resulting FoF halo largely conserves the mass from both progenitors. Thus, when using the total bound mass of the FoF as a halo mass proxy, the increment in the halo mass can have a good scaling with the increment in progenitor mass. On the other hand, the spherical overdensity mass definitions do not reflect the morphology and extent of the merging halos, so that the scaling between the progenitor mass and halo mass is more complex. This complexity can also be understood as an inconsistency between the boundary of the host halo and the orbital extent of its member particles and subhalos. Recent halo boundary definitions that incorporate the splashback or growth envelope of a halo, e.g., the depletion radius (Fong & Han, 2021; Gao et al., 2023), could potentially provide more self-consistent boundaries for the study of halo mergers and the resulting mass and subhalo distribution (Zhou & Han, 2023, 2025; Li & Han, 2021; Fong et al., 2022).
@@ 5.2 Correcting Excessive Mass from Unresolved Subhalos
Figure 11: The final subhalo mass functions in L600 at $z=0$ after being rescaled by $(M_{h}/10^{10}\mathrm{M}_{\odot}\;h^{-1})^{0.1}$. Dots of different colors show the results obatined from different halo mass bins. The errorbars show the uncertainty on the average final subhalo mass function estimated from the Poisson noise in the total subhalo count in each bin. The gray solid line indicates the best fit (Equation (18)).
Figure 12: Left: Comparisons of $g_{1}$ obtained from the Kanli CDM simulation and L600 without peak mass correction. All of them have been normalized by the best-fitting formula in Equation (9). The thick solid and light dashed lines represent results from L600 and Kanli CDM, respectively. In the plots, the PMFs are measured from two distinct host mass bins, $[10^{12},10^{13}]{\rm M_{\odot}}\,h^{-1}$ (blue points) and $[10^{13},10^{14}]{\rm M_{\odot}}\,h^{-1}$ (red points), in each simulation. Shaded regions indicate deviations of 5% from the fitting formula. Right: The comparisons of $g_{1}$ measured from two simulations after applying the mass correction to L600.
@@ 6 Conclusion
In this work, we have carried out a detailed analysis on the hierarchical origin of dark matter subhalos, using a set of N-body simulations. Applying the hbt+subhalo finder to the simulations, we classify subhalos into different levels according to their merger histories, and use the peak mass function (PMF) of each level to summarize their hierarchical origins, in which the peak mass of a subhalo is used as a proxy for the progenitor mass. Our main findings are summarized below.
  • The PMF is very close to a universal function across halo mass and redshift. A single double-Schechter fit can describe the level-1 subhalo PMF to $10\%$ accuracy down to a peak-to-host mass ratio of $\mu=5\times 10^{-4}$, for the halo mass range of $10^{11}\mathrm{M}_{\odot}\;h^{-1}-10^{15}\mathrm{M}_{\odot}\;h^{-1}$ and redshift range of $z=0-3$ covered in our analysis.
@@
Appendix A Halo-to-halo variation in the PMF
Figure 13: The ratio between the standard deviation of the PMFs among different halos and the Poission noise in the PMF of an average halo. Different colors show the results in different halo mass bins as labeled. The dashed line represents $\sigma/\sigma_{\rm Poisson}=1$.
To investigate the significance of halo-to-halo variation in the PMFs, we compare the standard deviation of subhalo counts ($\sigma=\sqrt{\sum_{i=1}^{N_{\rm h}}\frac{(N_{\rm sub,i}-\bar{N}_{\rm sub})^{2}}{N_{\rm h}}}$) among different halos and the Poisson noise ($\sigma_{\rm Poisson}=\sqrt{\bar{N}_{sub}}$).
In Figure 13, we plot the ratio of the measured scatter in the level-1 PMFs to the Poisson expectation, $\sigma/\sigma_{\mathrm{Poisson}}$, as a function of mass ratio $\mu$. For low mass ratios ($\mu\lesssim 10^{-2}$), the scatter is super-Poissonian ($\sigma/\sigma_{\mathrm{Poisson}}>1$), indicating a greater variability than that predicted by a Poisson distribution. In the intermediate range ($10^{-2}\leq\mu\leq 2\times 10^{-1}$), the scatter is sub-Poissonian ($\sigma/\sigma_{\mathrm{Poisson}}<1$). For higher mass ratios ($\mu>0.2$), the scatter aligns closely with the Poisson expectation ($\sigma/\sigma_{\mathrm{Poisson}}\approx 1$). These halo-to-halo variations are consistent with previous studies (e.g. Boylan-Kolchin et al., 2010; Jiang & van den Bosch, 2017; Chua et al., 2017). Additionally, the PMF scatter exhibits a mild dependence on halo mass, with the sub-Poissonian behavior more pronounced in massive halos. Finally, across the range of peak mass ratio $\mu$ considered in the paper, the observed scatter does not exceed 1.6 times the Poisson error.
Appendix B The cosmology dependence of the PMF in the EPS theory
In this appendix, we provide a theoretical interpretation for the PMF dependence on different cosmological parameters with the extended Press–Schechter (EPS) formalism . We follow Dong et al. (2022) to calculate the specific merger rate using the EPS theory (their Equation (11)) and then convert the result to the PMF (their Equation (4)). We use the Eisenstein & Hu (1998) transfer function to calculate the linear power spectrum and adopt a top-hat filter in Fourier space to compute the variance of fluctuation on a given scale. The predictions for the PMFs under different cosmological parameters are presented in Figure 14.
In the left panel, the predicted PMFs are nearly universal for $\mu$ values greater than $10^{-3}$, regardless of variations in $\Omega_{M_{0}}$ and $\sigma_{8}$. The slopes of the predicted PMFs show slight differences below $10^{-3}$, where larger $\Omega_{M_{0}}$ values result in steeper slopes. This cosmology dependence also suggests a weak redshift dependence in a given cosmology due to the evolution of the matter density and fluctuation amplitude with redshift. Verifying this residual dependence requires higher-resolution simulations than those used in the current work.
Finally, we emphasize that the EPS predictions here serve as a qualitative reference rather than a precise quantitative comparison to simulations. The theoretical mass function is not calibrated against simulation results using the halo mass definition adopted in this work, but instead directly follows the formulation of Lacey & Cole (1993). Furthermore, the EPS framework implicitly assumes consistent mass definitions for halos and subhalos when calculating the peak mass ratio, whereas our analysis adopts different definitions for the two. These inconsistencies can result in deviations in both the amplitude and slope when comparing the predicted PMFs with simulation results.